# Benchmarking Unsupervised Online IDS for Masquerade Attacks in CAN

### v1

@@ 4.5 Dataset
We report results on the proposed benchmark in each of the masquerade attacks provided in the ROAD dataset. Note that each of the masquerade attack files in ROAD contains hundreds of time series coming from hundreds IDs each of them with up to a few dozen signals. Table I shows the specific files we used from ROAD. In particular, we used the following attacks in increasing level of detection difficulty: correlated signal, max speedometer, max engine coolant temperature, reverse light on, and reverse light off. In the correlated signal attack, the correlation of the four wheel speed values is manipulated by changing their individual values. In the max speedometer and max engine coolant attacks, the speedometer and coolant temperature values are changed to their maximum. In the reverse light attacks, the state of the reverse light is modified so that it does not match the current gear of the vehicle (i.e., reverse light is off when the vehicle is in reverse and the reverse light is on when the vehicle is in forward gear).
TABLE I: CAN captures used from the ROAD dataset [10].
Description # Files Used Duration (min)
Training Dynamometer Various Ambient 10 108.2
Road Various Ambient 2 70.6
Total 12 12 178.8
Testing Correlated Signal Fabrication Attack 3 1.3
Correlated Signal Masquerade Attack 3 1.3
Fuzzing Fabrication Attack 3 0.7
Max Engine Coolant Temp Fabrication Attack 1 0.4
Max Engine Coolant Temp Masquerade Attack 1 0.4
Max Speedometer Fabrication Attack 3 3.9
Max Speedometer Masquerade Attack 3 3.9
Reverse Light Off Fabrication Attack 3 2.1
Reverse Light Off Masquerade Attack 3 2.1
Reverse Light On Fabrication Attack 3 3.2
Reverse Light On Masquerade Attack 3 3.2
Accelerator Attack (In Drive) 2 2.7
Accelerator Attack (In Reverse) 2 3.2
Total 33 13 10.9
We used the longest benign capture (i.e., ambient-highway-street-driving-long) for learning the normal state of the system ($\approx$ 1 hour of data). We use all the available attack data in the masquerade category, i.e., three files in the correlated attack category ($\approx$ 1.3 minutes), three files in max speedometer attack category ($\approx$ 3.9 minutes), one file for the max engine coolant attack category ($\approx$ 0.4 minutes), three files for reverse light on attack category ($\approx$ 3.2 minutes), and three files for the reverse light off attack category ($\approx$ 2.1 minutes).
4.6 Evaluation Metrics
@@ 5 Results
First, we compute and contrast the detection performance (based on AUC-ROC) of each algorithm per attack category in the ROAD dataset. In doing so, we display heatmaps (Section 5.1). We hypothesize that each algorithm is better suited for detecting specific attacks in the ROAD dataset and that there is a range of $(\omega,\delta)$ combinations that maximizes AUC-ROC.
Second, we summarize evaluation metrics using summary statistics derived from the heatmaps (Section 5.2). Specifically, we report the mean ($\mu$), standard deviation ($\sigma$), median ($\eta$), minimum ($\min$), including the combination of window length and offset that produces it, and maximum ($\max$), including the combination of window length and offset that produces it. This analysis supports our hypothesis that there are better suited detection algorithms for each attack category. Overall, Moriano22 detection algorithm produces, on average, the best detection performance among attacks in the ROAD dataset. In contrast, Ganesan17 detection algorithm produces the worst detection performance among attacks in the ROAD dataset. Furthermore, we found that the easiest attacks to detect are the max_engine_coolant_attack and correlated_attack. Furthermore, we find that the max_speedometer_attack and light_on_attack are the hardest to detect.
Lastly, we summarize TTW as a proxy of performance (Section 5.3). In particular, we report the mean ($\mu$), standard deviation ($\sigma$), median ($\eta$), minimum ($\min$), including the combination of window length and offset that produces it, and maximum ($\max$), including the combination of window length and offset that produces it. This analysis reveals that more sophisticated algorithms (e.g., Ganesan17 and Moriano22) incur performance overhead for producing better detection evaluation metrics. Overall, Matrix Correlation Distribution produces, on overage, the quickest TTW among attacks in the ROAD dataset. In contrast, Moriano22 detection algorithm is at least 4×4$\times$4 × slower. It is worth noting that detection algorithms that run faster (i.e., Matrix Correlation Distribution and Matrix Correlation Correlation) perform even better than Ganesan17 (see Table III). We did not found significant differences on TTW per attack categories.
5.1 Evaluation Metrics Contrast
@@
5.2 Evaluation Metrics Summary
@@
Finally, Fig. 8 compares each detection method on the basis of summary statistics for each of attack type using dot plots. A desirable online algorithm for detecting masquerade attacks in CAN by processing streams of time series data should have high values of $\mu$, $\eta$, $\min$, and $\max$ and lower values of $\sigma$ in the range $[0,1]$. We notice that there is no free lunch in the detection of masquerade attacks in CAN in an online setting as there is no single algorithm that is optimal for detecting each attack type. Notably, Ganesan17, the worst performing algorithms across all attack types is the best performing detection algorithm for the hardest attack to detect, i.e., max_speedometer_attack.
@@
5.3 Performance Metrics Summary
Table III summarizes TTW results based on heatmaps that we do not display here. We place attack categories in the rows and attack detection methods in the columns. This means that each cell in this table shows the TTW of a specific detection method in a particular attack from the ROAD dataset. We focus on summary statistics including the mean ($\mu$), standard deviation ($\sigma$), median ($\eta$), minimum ($\min$), and maximum ($\max$). In addition, we also show the average and standard deviations from the minimum values for each attack detection methods and attack types in the ROAD dataset. They are displayed as the last row and column in the table.
@@ 5.3 Performance Metrics Summary
Finally, Fig. 9 compares each detection method on the basis of TTW summary statistics for each attack type using dot plots. A desirable online algorithm for detecting masquerade attacks in CAN by processing streams of time series data should have low values of $\mu$, $\eta$, $\min$, $\max$ and lower values of $\sigma$ in the range $[0,1]$. Remarkably, Matrix Correlation Distribution method is the quickest detection algorithm one order of magnitude lower than the others.
@@
TABLE III: Summary of evaluation metrics based on TTW in miliseconds. We show the mean ($\mu$), standard deviation ($\sigma$), median ($\eta$), minimum ($\min$) including the combination of window length and offset that produces it, and maximum ($\max$) including the combination of window length and offset that produces it. Attack types are placed in the rows while detection methods are placed in the columns. We also report averages and standard deviation of the minimum values obtained by each attack and detection method denoted by $\overline{min}$ and $\sigma_{min}$.
@@ 6 Discussion
Use of Default Parameter Values in Detection Algorithms: We used default parameters values as the original algorithms in Ganesan17 and Moriano22. Therefore we did not explore the use of hyperparameter tuning on these algorithms that could have an effect in their performance.

### v2

@@ IV-D Dataset
We used the longest benign capture in the ROAD dataset (i.e., ambient-highway-street-driving-long) for learning the normal state of the system ($\approx$ 1 hour of data). We report results on the proposed benchmark in each of the masquerade attacks provided in the ROAD dataset. Note that each of the masquerade attack files in ROAD contains hundreds of time series coming from hundreds IDs each of them with up to a few dozen signals. Table I shows the specific files we used from ROAD. In particular, we used the following attacks in increasing level of detection difficulty: correlated signal, max speedometer, max engine coolant temperature, reverse light on, and reverse light off. In the correlated signal attack, the correlation of the four wheel speed values is manipulated by changing their individual values. In the max speedometer and max engine coolant attacks, the speedometer and coolant temperature values are changed to their maximum. In the reverse light attacks, the state of the reverse light is modified so that it does not match the current gear of the vehicle (i.e., reverse light is off when the vehicle is in reverse and the reverse light is on when the vehicle is in forward gear).
TABLE I: Masquerade attack characteristics in the ROAD dataset [10].
[…]
IV-E Evaluation Metrics
@@ V Results
First, we compute and contrast the detection performance (based on AUC-ROC) of each algorithm per attack category in the ROAD dataset. In doing so, we display heatmaps (Section V-A). We hypothesize that each algorithm is better suited for detecting specific attacks in the ROAD dataset and that there is a range of $(\omega,\delta)$ combinations that maximizes AUC-ROC.
Second, we summarize evaluation metrics using summary statistics derived from the heatmaps (Section V-B). Specifically, we report the mean ($\mu$), standard deviation ($\sigma$), median ($\eta$), minimum ($\min$), including the combination of window length and offset that produces it, and maximum ($\max$), including the combination of window length and offset that produces it. This analysis supports our hypothesis that there are better suited detection algorithms for each attack category. Overall, Moriano22 detection algorithm produces, on average, the best detection performance among attacks in the ROAD dataset. In contrast, Ganesan17 detection algorithm produces the worst detection performance among attacks in the ROAD dataset. Furthermore, we found that the easiest attacks to detect are the max_engine_coolant_attack and correlated_attack. Furthermore, we find that the max_speedometer_attack and light_on_attack are the hardest to detect. We also optimize the hyperparameters for the best-performing method, i.e., Moriano22. Using grid search, we demonstrate the potential for AUC-ROC improvement through hyperparameter tuning.
Lastly, we summarize TTW as a proxy of performance (Section V-C). In particular, we report the mean ($\mu$), standard deviation ($\sigma$), median ($\eta$), minimum ($\min$), including the combination of window length and offset that produces it, and maximum ($\max$), including the combination of window length and offset that produces it. This analysis reveals that more sophisticated algorithms (e.g., Ganesan17 and Moriano22) incur performance overhead for producing better detection evaluation metrics. Overall, Matrix Correlation Distribution produces, on overage, the quickest TTW among attacks in the ROAD dataset. In contrast, the Moriano22 detection algorithm runs at least four times slower, as it uses a standard AHC implementation rather than a faster approximate alternative [65]. It is worth noting that detection algorithms that run faster (i.e., Matrix Correlation Distribution and Matrix Correlation Correlation) perform even better than Ganesan17 (see Table IV). We did not found significant differences on TTW per attack categories.
V-A Evaluation Metrics Contrast
@@
Masquerade attacks in the ROAD dataset are sustained and stealthy, lasting about 17–72 seconds (see Table I). We tuned the sliding window parameters ($\omega$ and $\delta$) to match attack duration while keeping computation manageable. For example, a window size of 400 samples gives enough temporal context to detect slow-developing anomalies. An offset of 380 samples allows overlap between windows, which helps avoid missed detections at window edges. This offset does not delay response but balances detection coverage and computational cost. Smaller offsets (e.g., 100) increase load, while larger ones (e.g., 400) may miss attacks. These parameters match the timing of real attacks. Note that masquerade attacks do not show up in single CAN frames. They imitate normal traffic in short bursts but become apparent as gradual deviations. As shown by Jedh et al. [66], detecting these attacks requires looking across multiple frames, not just isolated ones.
@@
V-B Evaluation Metrics Summary
@@
Additionally, Fig. 8 compares each detection method on the basis of summary statistics for each of attack type using dot plots. A desirable online algorithm for detecting masquerade attacks in CAN by processing streams of time series data should have high values of $\mu$, $\eta$, $\min$, and $\max$ and lower values of $\sigma$ in the range $[0,1]$. We notice that there is no free lunch in the detection of masquerade attacks in CAN in an online setting as there is no single algorithm that is optimal for detecting each attack type. Notably, Ganesan17, the worst performing algorithms across all attack types is the best performing detection algorithm for the hardest attack to detect, i.e., max_speedometer_attack.
@@
Finally, we performed grid search over predefined hyperparameter ranges for Moriano22. Specifically, we varied $r\in\{-5,-3,\ldots,3\}$ and $\alpha\in\{0.1,0.2,\dots,0.8\}$. These ranges cover a variety of standard configurations. To ensure a thorough evaluation, we applied this procedure to the specific combination of $\omega$ and $\delta$, where Moriano22 performed the worst among all attack scenarios. This allowed us to analyze the worst-case scenario and assess the performance improvements achievable by tuning hyperparameters. We visualized the sensitivity of AUC-ROC scores to these parameters using heatmaps. For each attack category, we computed the average performance and summarized it in a single heatmap. This approach helps interpret the effect of hyperparameter tuning on the method’s performance. Figure 9 illustrates how AUC-ROC scores improve with optimized combinations of $r$ and $\alpha$. Notably, scores increased significantly from their minimum values (see Table 2 column Moriano22) to their maximum values shown in the heatmaps, except for the light_off_attack category. In addition, Table III compares the AUC-ROC values of the Moriano22 method before and after hyperparameter optimization. This analysis confirms the robustness of the proposed method under optimized settings while highlighting areas for further optimization.
Figure 9: AUC-ROC values for Moriano22 across different combinations of $r$ and $\alpha$ in the ROAD dataset attacks: (a) correlated attack, (b) max engine coolant temperature attack, (c) max speedometer attack, (d) reverse light off attack, and (e) reverse light on attack.
TABLE III: Comparison of AUC-ROC values for the Moriano22 method before and after hyperparameter optimization. The “before” column shows values from Table 2, where the method produces the lowest result. The “after” column presents values obtained using the window length and offset combination from the “before” column. The corresponding $r$ and $\alpha$ combinations that achieve the maximum values are shown in parentheses. Attack types are listed in the rows, while results before and after optimization are in the columns. The last column reports the absolute percentage change.
Attack/Method Moriano22 Before Moriano22 After Absolute
Optimization Optimization Change (%)
correlated_attack 0.15 0.84 (3, 0.8) +69%
max_engine_attack 0.00 0.10 (3, 0.3) +10%
max_speedometer_attack 0.42 0.57 (-5, 0.1) +15%
light_off_attack 0.44 0.40 (-5, 0.1) -4%
light_on_attack 0.51 0.66 (-5, 0.1) 15%
V-C Performance Metrics Summary
Table IV summarizes TTW results based on heatmaps that we do not display here. We place attack categories in the rows and attack detection methods in the columns. This means that each cell in this table shows the TTW of a specific detection method in a particular attack from the ROAD dataset. We focus on summary statistics including the mean ($\mu$), standard deviation ($\sigma$), median ($\eta$), minimum ($\min$), and maximum ($\max$). In addition, we also show the average and standard deviations from the minimum values for each attack detection methods and attack types in the ROAD dataset. They are displayed as the last row and column in the table.
@@ V-C Performance Metrics Summary
Finally, Fig. 10 compares each detection method on the basis of TTW summary statistics for each attack type using dot plots. A desirable online algorithm for detecting masquerade attacks in CAN by processing streams of time series data should have low values of $\mu$, $\eta$, $\min$, $\max$ and lower values of $\sigma$ in the range $[0,1]$. Remarkably, Matrix Correlation Distribution method is the quickest detection algorithm one order of magnitude lower than the others. Note that, the benchmarked algorithms infer in near real-time, with TTW values typically within a few milliseconds. Since human driver response times range from 0.7 to 1.5 seconds [67], the algorithms meet the requirement of inferring well below the lower bound of 0.7 seconds.
@@
TABLE IV: Summary of performance metrics based on TTW in miliseconds. We show the mean ($\mu$), standard deviation ($\sigma$), median ($\eta$), minimum ($\min$) including the combination of window length and offset that produces it, and maximum ($\max$) including the combination of window length and offset that produces it. Attack types are placed in the rows while detection methods are placed in the columns. We also report averages and standard deviation of the minimum values obtained by each attack and detection method denoted by $\overline{min}$ and $\sigma_{min}$.
@@ VI Discussion
Although Moriano22 outperformed the other methods, achieving the best detection results across attack types, it showed lower average AUC-ROC values for the correlated_attack (i.e., $\mu$ = 0.35). This result likely stems from the windowing approach used in the testing, which limits capturing longer-term effects of masquerade attacks. These longer-term effects may hinder the formation of consistent cluster hierarchies and could be further explored with other time series prequential evaluation methods [47]. However, Moriano22 outperformed the other methods on the remaining attacks—max_engine_attack, max_speedometer_attack, light_off_attack, and light_on_attack. This indicates greater robustness in detecting masquerade attacks, particularly in the presence of subtle data manipulations such as those affecting the engine coolant, speedometer, and reverse light signals. Given the variability in performance across attacks, we expect that combining the benchmarked algorithms through ensemble learning could improve overall robustness. An ensemble could exploit the strengths of individual methods and compensate for their weaknesses, especially when detecting subtle or sustained attacks.
@@
Use of Default Parameter Values in Detection Algorithms: We used default parameters values as the original algorithms in Ganesan17 and Moriano22. We optimized the hyperparameters for Moriano22, focusing on a single combination of $\omega$ and $\delta$ that achieved the lowest AUC-ROC with default settings. This approach confirms the method’s robustness under optimized conditions and identifies areas for further improvement.